Format your numerical answers to 4 decimals.

  1. A survey of 200 college graduates who have been working for at least three years found that 60 owned only mutual funds, 40 owned only stocks, 40 owned only bonds and 50 owned all three stocks, bonds and mutual funds.
    1. What is the probability that an individual owns a stock?
    2. What is the probability that an individual owns a bond?
    3. What is the probability that an individual owns a mutual fund?
    4. What is the probability that an individual owns neither stock nor bond nor mutual fund?
    5. What is the probability that an individual owns either a stock or a bond or a mutual fund?
  1. In determining automobile-mileage ratings, it was found that the mpg (X) for a certain model is normally distributed, with mean of 32 mpg and a standard deviation of 2.4 mpg. Find the following:
    1. P ( X < 30)
    2. P ( 28 < X < 31)
    3. P ( X > 33)
    4. P ( X > 35)
    5. The mileage rating that the upper 1% of cars achieve.
Answer #1

Solution: -

Total college graduates = 200

(I) Total individual that have a stock = 40 +50 = 90

Now the probability that individual owns a stock = 90/200 = 0.45

(II) Total individuals that owns a bond = 40 +50 = 90

Now the probability that individual owns a bond = 90/200

= 0.45

(III) Total individuals that owns a mutual fund = 60 + 50 = 110

Now the probability that individuals own a mutual = 110/200

= 0.55

(IV) Now the persons that have neither stock nor bond nor mutual fund = (200 - 60 - 40 - 40 -50) = 10

Now their probabilty = 10/200 = 0.05

As per HomeworkLib policy we need to answer maximum four subparts in a question so please ask other parts as seperate one.... Sorry